# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ II System Model
$\xi_{i}\triangleq{Q^{-1}(\varsigma_{i})}/{\text{ln}(2)\sqrt{N_{t}}}$, $Q^{-1}(\cdot)$ is the inverse of the Gaussian Q-function, $N_{t}\triangleq\mathcal{B}t_{t}$ is the packet length, $\mathcal{B}$ is the bandwidth, $t_{t}$ is the transmission duration, $\varsigma_{k}$ is the decoding error probability[26], $\varsigma_{e}$ is the information leakage[26], and $V_{k}$ and $V_{e}$ are the dispersion factors defined as [26],
$V_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\rho_{i}}/{\upsilon_{i}})\right),i\in\{k,e\},$ (11)
$\widehat{V}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\widehat{\rho}_{i}}/{\widehat{\upsilon}_{i}})\right),i\in\{k,e\},$ (12)
@@
Unlike the traditional SR definition in the LBR, the FBR induces more constraints on the SR. Specifically, the reliable transmission in FBR requires that the decoding error probability $\varsigma_{k}$ at user-$k$ is not larger than the maximum decoding error probability in FBR $\varsigma_{\max}$. Additionally, the information leakage constraint imposes that the information leakage $\varsigma_{e}$ does not exceed $\varsigma_{\max}$, the maximum information leakage in FBR [26]. These two conditions distinguish the FBR case from its LBR counterpart. One may notice that when the transmission duration $t_{t}$ approaches infinity, the dispersion factor $V_{i}$ reaches zero, and the FBR definitions (7) and (8) reduce to the traditional LBR-SR case, which can be expressed as [27]:
@@ III MAXMIN SR under perfect CSI in FBR systems
The optimization problem ($\mathcal{P}1$) is non-convex since the objective function (15a) is not concave and the UMC (15c) is non-convex. To tackle this non-convex problem, one can employ the AO technique [1]. Specifically, at iteration $(\iota)$, the feasible point $(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$ is generated from ($\mathcal{P}1$) by solving two sub-problems, the first one is to optimize $\boldsymbol{w}$ with a fixed $\boldsymbol{\theta}$:
@@
$C_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ $\geq q_{1,k_{e}}^{(\iota)}+n_{1,k_{e}}^{(\iota)}+2\Re\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\left\{\left\langle\boldsymbol{m}_{j,k_{e}}^{(\iota)},\boldsymbol{w}_{j}\right\rangle\right\}-\frac{\rho_{e}^{(\iota)}}{1+\rho_{e}^{(\iota)}}\left(\sum\limits_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}|^{2}\right)-\frac{1}{\upsilon_{e}^{(\iota)}}\sum\limits_{j=1}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}|^{2},$ (20)
${C_{k}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ $\geq{q}_{1,k}^{(\iota+1)}+2\Re\left\{{\sum}_{n=1}^{N}{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)e^{j\theta_{n}}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k}^{(\iota+1)}e^{j\boldsymbol{\theta}},$ (24)
$C_{e}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ $\geq q_{1,k_{e}}^{(\iota+1)}+n_{1,k_{e}}^{(\iota+1)}+2\mathfrak{R}\left\{{\sum}_{n=1}^{N}{\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)e^{(j\theta_{n})}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k_{e}}^{(\iota+1)}e^{j\boldsymbol{\theta}}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{k_{e},j}^{(\iota+1)}e^{j\boldsymbol{\theta}}.$ (25)
@@ III-1 Sub-Problem for Optimizing the Beamforming Vectors
We fix $\boldsymbol{\theta}$ given $\boldsymbol{w}^{(\iota)}$ and solve the problem ($\mathcal{P}1$) to obtain $\boldsymbol{w}^{(\iota+1)}$ satisfying ${\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})\geq{\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})$. We start by linearizing the objective function in (15a), which consists of four parts: the user-$k$’s SR $C_{k}(\boldsymbol{w},\boldsymbol{\theta})$, Eve’s negative eavesdropping rate $C_{e}(\boldsymbol{w},\boldsymbol{\theta})$, the user-$k$ dispersion factor $V_{k}(\boldsymbol{w},\boldsymbol{\theta})$, and Eve’s dispersion factor $V_{e}(\boldsymbol{w},\boldsymbol{\theta})$ as shown in (7).
First, we adopt the idea in [28] to convert user-$k$’s SR into a linear form. Specifically, using the inequality (74) in Appendix D, let’s define $\boldsymbol{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{k}$, ϝ$\boldsymbol{\digamma}\triangleq\rho_{k}$, $\hat{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{k}^{(\iota)}$ and ϝ^$\hat{\digamma}\triangleq\rho_{k}$, hence, the users’ data rate can be written as:
$C_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})\geq q_{1,k}^{(\iota)}+2\Re\left\{\left\langle\boldsymbol{m}_{k,k}^{(\iota)},\boldsymbol{w}_{k}\right\rangle\right\}-n_{1,k}^{(\iota)}\upsilon_{{}_{k}}^{(\iota)},$ (18)
@@
$q_{1,k}^{(\iota)}$ $\triangleq C_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})-\gamma_{k}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})-\sigma_{k}n_{1,k}^{(\iota)},$
$\boldsymbol{m}^{(\iota)}_{k,k}$ $\triangleq{\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{{}_{k}}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}^{(\iota)}_{k}}/{\rho_{k}^{(\iota)}},$
$n_{1,k}^{(\iota)}$ $\triangleq{1}/{\rho_{k}^{(\iota)}}-{1}/{\upsilon_{{}_{k}}^{(\iota)}}.$
@@
$-\ln(1+\gamma_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)}))$ $\triangleq\overset{a_{1}}{\overbrace{\ln(1+\rho_{e}^{(\iota)})}}-\overset{a_{2}}{\overbrace{\ln(1+\upsilon_{e}^{(\iota)})}}.$ (19)
@@
$q_{1,k_{e}}^{(\iota)}$ $\triangleq\ln\left(\rho_{e}^{(\iota)}\right)-\rho_{e}^{(\iota)}-\ln\left(\upsilon_{e}^{(\iota)}\right)+1,$
$n_{1,k_{e}}^{(\iota)}$ $\triangleq\left(-\frac{\rho_{e}^{(\iota)}}{1+\rho_{e}^{(\iota)}}-\frac{1}{\upsilon_{e}^{(\iota)}}\right)\sigma_{e},$
$\boldsymbol{m}_{k_{e},j}^{(\iota)}$ $\triangleq\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{w}_{j}^{(\iota)}.$
Next, we linearize the FBR dispersion factors $V_{i}(\boldsymbol{w},\boldsymbol{\theta}),i\in\{k,e\}$. Using (11), the inequalities (72) and (73) in Appendix D, and by defining $x\triangleq V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta})$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\varsigma_{i}$, the dispersion factors can be expressed as:
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})}\leq q_{2,i}^{(\iota)}-2\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\Re\left\{\left\langle\boldsymbol{d}_{j,i}^{(\iota)},\boldsymbol{w}_{j}\right\rangle\right\}-n_{2,i}^{(\iota)}\upsilon^{(\iota)},$ (21)
@@
$q_{2,i}^{(\iota)}$ $\triangleq\xi_{i}\left(\frac{\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}}{2}+\frac{\left(\upsilon^{(\iota)}_{i}\right)^{2}+\rho^{(\iota)}_{i}\sigma_{i}-2\upsilon^{(\iota)}_{i}\sigma_{i}}{\left(\upsilon^{(\iota)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}}\right),$
$\boldsymbol{d}_{j,i}^{(\iota)}$ $\triangleq\frac{\xi_{i}[\boldsymbol{h}_{i}(\boldsymbol{\theta}^{(\iota)})]^{2}\boldsymbol{w}^{(\iota)}_{j}}{\upsilon^{(\iota)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}},i\in\{k,e\},$
$n_{2,i}^{(\iota)}$ $\triangleq\frac{\xi_{i}\rho^{(\iota)}_{i}}{\left(\upsilon^{(\iota)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}},i\in\{k,e\}.$
By substitution (18), (20), and (21) into (7), the FBR-SR can be written as:
$\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)})$ $\geq{q}_{k}^{(\iota)}+2\Re\left\{\left\langle{\boldsymbol{m}}_{k}^{(\iota)},\boldsymbol{w}_{k}\right\rangle\right\}-(\boldsymbol{w}_{k})^{H}{\boldsymbol{\psi}}^{(\iota)}_{k}\boldsymbol{w}_{k},$ (22)
where
${q}_{k}^{(\iota)}$ $\triangleq q_{1,k}^{(\iota)}+q_{1,k_{e}}^{(\iota)}+n_{1,k_{e}}^{(\iota)}+q_{2,k}^{(\iota)}+q_{2,k_{e}}^{(\iota)},$
${\boldsymbol{m}}_{k}^{(\iota)}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{m}_{j,k}^{(\iota)}+{\sum}_{j=1,j\neq k}^{\mathcal{K}}\left(\boldsymbol{m}_{k_{e},j}^{(\iota)}-\boldsymbol{d}_{j,k}^{(\iota)}-\boldsymbol{d}_{j,k_{e}}^{(\iota)}\right),$
$\boldsymbol{\psi}^{(\iota)}_{k}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}n_{j}^{(\iota)}\boldsymbol{h}_{j}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{j}(\boldsymbol{\theta}^{(\iota)})+n_{k_{e}}^{(\iota)}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(\iota)}),$
$n_{k}^{(\iota)}$ $\triangleq n_{1,k}^{(\iota)}+n_{2,k}^{(\iota)},$
$n_{k_{e}}^{(\iota)}$ $\triangleq\frac{\rho_{e}^{(\iota)}}{\left(1+\rho_{e}^{(\iota)}\right)}+\frac{1}{\upsilon_{e}^{(\iota)}}+n_{2,k_{e}}.$
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})}$ $\leq q_{2,i}^{(\iota+1)}-2\Re\left\{{\sum}_{n=1}^{N}\boldsymbol{d}_{i}^{(\iota+1)}e^{(j\theta_{n})}\right\}+\left(e^{(j\boldsymbol{\theta})}\right)^{H}\boldsymbol{\varphi}_{2,i}^{(\iota+1)}e^{(j\boldsymbol{\theta})},i\in\{k,e\}.$ (26)
$q_{2,i}^{(\iota+1)}$ $\triangleq\frac{\xi_{i}{\left(\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}+2\left(\upsilon^{(\iota+1)}_{i}\right)^{2}+2\sigma_{i}\rho_{i}^{(\iota+1)}-4\sigma_{i}\left(\upsilon^{(\iota+1)}_{i}\right)\right)}}{{\left(2\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}},$ (27)
@@
$\boldsymbol{d}_{i,j}^{(\iota+1)}(n)$ $\triangleq{\xi_{i}}\bigg{/}{\left(\upsilon^{(\iota+1)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}\left(\boldsymbol{w}^{(\iota+1)}_{j}\right)^{H}\boldsymbol{h}_{i}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{u}_{i}\Xi_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{j},n\in N,j\in K,j\neq k,$ (28)
@@
$q_{k}^{(\iota+1)}$ $\triangleq{q}_{1,k}^{(\iota+1)}+{q}_{1,k_{e}}^{(\iota+1)}+q_{2,k}^{(\iota+1)}+q_{2,k_{e}}^{(\iota+1)}+n_{1,ke}^{(\iota+1)}-\left(e^{j\boldsymbol{\theta}^{(\iota)}}\right)^{H}\boldsymbol{\varphi}_{k}^{(\iota+1)}e^{j\boldsymbol{\theta}}-2\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(\iota+1)}\right)N,$ (30)
@@
${\boldsymbol{m}}^{(\iota+1)}_{k}(n)$ $\triangleq{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)+{\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)+{\boldsymbol{d}}^{(\iota+1)}_{k}(n)+{\boldsymbol{d}}^{(\iota+1)}_{k_{e}}(n)+{\sum}_{m=1}^{N}e^{-j\theta_{m}^{(\iota)}}\boldsymbol{\varphi}_{k}^{(\iota+1)}(m,n)+\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(\iota+1)}\right).$ (31)

### v3

@@ II System Model
$\xi_{i}\triangleq{Q^{-1}(\tau_{i})}/{\text{ln}(2)\sqrt{N_{t}}}$, $Q^{-1}(\cdot)$ is the inverse of the Gaussian Q-function, $N_{t}\triangleq\mathcal{B}t_{t}$ is the packet length, $\mathcal{B}$ is the bandwidth, $t_{t}$ is the transmission duration, $\tau_{k}$ is the decoding error probability, $\tau_{e}$ is the information leakage[27], and $V_{k}$ and $V_{e}$ are the dispersion factors defined as [27],
$V_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\gamma_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\rho_{i}}/{\upsilon_{i}})\right),$ (11)
$\widehat{V}_{i}(\boldsymbol{w},\boldsymbol{\theta})$ $\triangleq\frac{2\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta})}{(1+\widehat{\gamma}_{i}(\boldsymbol{w},\boldsymbol{\theta}))}\triangleq 2\left(1-({\widehat{\rho}_{i}}/{\widehat{\upsilon}_{i}})\right),$ (12)
@@
Unlike the traditional SR definition in the LBR, the FBR imposes more constraints on the SR. Specifically, the reliable transmission in FBR requires that the decoding error probability $\tau_{k}$ at user-$k$ is not larger than the maximum decoding error probability in FBR $\tau_{\max}$, the information leakage constraint imposes that the information leakage $\tau_{e}$ does not exceed $\tau_{e_{\max}}$, the maximum information leakage in FBR [27]. Additionally, the transmission duration $t_{t}$ doesn’t exceed the maximum transmission duration $t_{\max}$. Moreover, ensuring secure communication in the FBR regime with packet length $N_{t}$, is more challenging than in its LBR counterpart. This increased difficulty arises from the inherent limitations on the achievable SR in the FBR setting, which are due to the combined effects of non-negligible decoding error probabilities and information leakage. These two conditions distinguish the FBR case from its LBR counterpart. One may notice that when the transmission duration $t_{t}$ approaches infinity, the dispersion factor $V_{i}$ reaches zero, and the FBR definitions (7) and (8) reduce to the traditional LBR-SR case, which can be expressed as [28]:
@@ III MAXMIN SR under perfect CSI in FBR systems
The optimization problem ($\mathcal{P}1$) is nonconvex since the objective function (15a) is not concave the UMC (15c) is nonconvex. To tackle this nonconvex problem, one can employ the AO technique [1]. Specifically, at iteration $(n)$, the feasible point $(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})$ is generated from ($\mathcal{P}1$) by solving two sub-problems, the first one is to optimize $\boldsymbol{w}$ with a fixed $\boldsymbol{\theta}$:
@@
$C_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(n)})$ $\geq x_{1,k_{e}}^{(n)}+z_{1,k_{e}}^{(n)}+2\Re\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\left\{\left\langle\boldsymbol{y}_{j,k_{e}}^{(n)},\boldsymbol{w}_{j}\right\rangle\right\}-\frac{\rho_{e}^{(n)}}{1+\rho_{e}^{(n)}}\left(\sum\limits_{j=1,j\neq k}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}_{j}|^{2}\right)-\frac{1}{\upsilon_{e}^{(n)}}\sum\limits_{j=1}^{\mathcal{K}}|\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}_{j}|^{2},$ (20)
${C_{k}}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta})$ $\geq{x}_{1,k}^{(n+1)}+2\Re\left\{{\sum}_{m=1}^{M}{\boldsymbol{y}}^{(n+1)}_{k,k}(m)e^{j\theta_{m}}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k}^{(n+1)}e^{j\boldsymbol{\theta}},$ (24)
$C_{e}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta})$ $\geq x_{1,k_{e}}^{(n+1)}+z_{1,k_{e}}^{(n+1)}+2\mathfrak{R}\left\{{\sum}_{m=1}^{M}{\boldsymbol{y}}^{(n+1)}_{k_{e}}(m)e^{(j\theta_{m})}\right\}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{1,k_{e}}^{(n+1)}e^{j\boldsymbol{\theta}}+\left(e^{j\boldsymbol{\theta}}\right)^{H}{\boldsymbol{\varphi}}_{k_{e},j}^{(n+1)}e^{j\boldsymbol{\theta}}.$ (25)
@@ III-1 Sub-Problem for Optimizing the Beamforming Vectors
We fix $\boldsymbol{\theta}$ given $\boldsymbol{w}^{(n)}$ and solve the problem ($\mathcal{P}1$) to obtain $\boldsymbol{w}^{(n+1)}$ satisfying ${\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})\geq{\mathcal{S}}^{\mathcal{F}}_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})$. We start by linearizing the objective function in (15a), which consists of four parts: the user-$k$’s SR $C_{k}(\boldsymbol{w},\boldsymbol{\theta})$, Eve’s negative eavesdropping rate $C_{e}(\boldsymbol{w},\boldsymbol{\theta})$, the user-$k$ dispersion factor $V_{k}(\boldsymbol{w},\boldsymbol{\theta})$, and Eve’s dispersion factor $V_{e}(\boldsymbol{w},\boldsymbol{\theta})$ as shown in (7).
First, we adopt the idea in [29] to convert user-$k$’s SR into a linear form. Specifically, using the inequality (71) in Appendix D, let’s define $\boldsymbol{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}_{k}$, $\boldsymbol{\digamma}\triangleq\rho_{k}$, $\hat{\Lambda}\triangleq\boldsymbol{h}_{k}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}_{k}^{(n)}$ and $\hat{\digamma}\triangleq\rho_{k}$, hence, the users’ data rate can be written as:
$C_{k}(\boldsymbol{w},\boldsymbol{\theta}^{(n)})\geq x_{1,k}^{(n)}+2\Re\left\{\left\langle\boldsymbol{y}_{k,k}^{(n)},\boldsymbol{w}_{k}\right\rangle\right\}-z_{1,k}^{(n)}\upsilon_{{}_{k}}^{(n)},$ (18)
@@
$x_{1,k}^{(n)}$ $\triangleq C_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})-\gamma_{k}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})-\sigma_{k}z_{1,k}^{(n)},$
$\boldsymbol{y}^{(n)}_{k,k}$ $\triangleq{\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{h}_{{}_{k}}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}^{(n)}_{k}}/{\rho_{k}^{(n)}},$
$z_{1,k}^{(n)}$ $\triangleq{1}/{\rho_{k}^{(n)}}-{1}/{\upsilon_{{}_{k}}^{(n)}}.$
@@
$-\ln(1+\gamma_{e}(\boldsymbol{w},\boldsymbol{\theta}^{(n)}))$ $\triangleq\overset{a_{1}}{\overbrace{\ln(1+\rho_{e}^{(n)})}}-\overset{a_{2}}{\overbrace{\ln(1+\upsilon_{e}^{(n)})}}.$ (19)
@@
$x_{1,k_{e}}^{(n)}$ $\triangleq\ln\left(\rho_{e}^{(n)}\right)-\rho_{e}^{(n)}-\ln\left(\upsilon_{e}^{(n)}\right)+1,$
$z_{1,k_{e}}^{(n)}$ $\triangleq\left(-\dfrac{\rho_{e}^{(n)}}{1+\rho_{e}^{(n)}}-\dfrac{1}{\upsilon_{e}^{(n)}}\right)\sigma_{e},$
$\boldsymbol{y}_{k_{e},j}^{(n)}$ $\triangleq\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(n)})\boldsymbol{w}_{j}^{(n)}.$
Next, we linearize the FBR dispersion factors $V_{i}(\boldsymbol{w},\boldsymbol{\theta}),i\in\{k,e\}$. Using (11), the inequalities (69) and (70) in Appendix D, and by defining $x\triangleq V_{i}(\boldsymbol{w}^{(n)},\boldsymbol{\theta})$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\tau_{i}$, the dispersion factors can be expressed as:
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w},\boldsymbol{\theta}^{(n)})}\leq x_{2,i}^{(n)}-2\sum\limits_{j=1,j\neq k}^{\mathcal{K}}\Re\left\{\left\langle\boldsymbol{f}_{j,i}^{(n)},\boldsymbol{w}_{j}\right\rangle\right\}-z_{2,i}^{(n)}\upsilon^{(n)},$ (21)
where,
$x_{2,i}^{(n)}$ $\triangleq\xi_{i}\left(\frac{\sqrt{V_{i}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}}{2}+\frac{\left(\upsilon^{(n)}_{i}\right)^{2}+\rho^{(n)}_{i}\sigma_{i}-2\upsilon^{(n)}_{i}\sigma_{i}}{\left(\upsilon^{(n)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}}\right),$
$\boldsymbol{f}_{j,i}^{(n)}$ $\triangleq{\xi_{i}[\boldsymbol{h}_{i}(\boldsymbol{\theta}^{(n)})]^{2}\boldsymbol{w}^{(n)}_{j}}/{\upsilon^{(n)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}},$
$z_{2,i}^{(n)}$ $\triangleq\dfrac{\xi_{i}\rho^{(n)}_{i}}{\left(\upsilon^{(n)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}}.$
By substitution (18), (20), and (21) into (7), the approximated surrogate FBR-SR function can be written as:
$\widetilde{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(n)})$ $\geq{x}_{k}^{(n)}+2\Re\left\{\left\langle{\boldsymbol{y}}_{k}^{(n)},\boldsymbol{w}_{k}\right\rangle\right\}-(\boldsymbol{w}_{k})^{H}{\boldsymbol{\psi}}^{(n)}_{k}\boldsymbol{w}_{k},$ (22)
@@
${x}_{k}^{(n)}$ $\triangleq x_{1,k}^{(n)}+x_{1,k_{e}}^{(n)}+z_{1,k_{e}}^{(n)}+x_{2,k}^{(n)}+x_{2,k_{e}}^{(n)},$
${\boldsymbol{y}}_{k}^{(n)}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{y}_{j,k}^{(n)}+{\sum}_{j=1,j\neq k}^{\mathcal{K}}\left(\boldsymbol{y}_{k_{e},j}^{(n)}-\boldsymbol{f}_{j,k}^{(n)}-\boldsymbol{f}_{j,k_{e}}^{(n)}\right),$
$\boldsymbol{\psi}^{(n)}_{k}$ $\triangleq{\sum}_{j=1}^{\mathcal{K}}z_{j}^{(n)}\boldsymbol{h}_{j}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{h}_{j}(\boldsymbol{\theta}^{(n)})+z_{k_{e}}^{(n)}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{h}_{e}(\boldsymbol{\theta}^{(n)}),$
$z_{k}^{(n)}$ $\triangleq z_{1,k}^{(n)}+z_{2,k}^{(n)},$
$z_{k_{e}}^{(n)}$ $\triangleq{\rho_{e}^{(n)}}/{\left(1+\rho_{e}^{(n)}\right)}+{1}/{\upsilon_{e}^{(n)}}+z_{2,k_{e}}.$
$\xi_{i}\sqrt{V_{i}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta})}$ $\leq x_{2,i}^{(n+1)}-2\Re\left\{{\sum}_{m=1}^{M}\boldsymbol{f}_{i}^{(n+1)}e^{(j\theta_{m})}\right\}+\left(e^{(j\boldsymbol{\theta})}\right)^{H}\boldsymbol{\varphi}_{2,i}^{(n+1)}e^{(j\boldsymbol{\theta})},i\in\{k,e\}.$ (26)
@@
$x_{2,i}^{(n+1)}$ $\triangleq\frac{\xi_{i}{\left(\left(\upsilon^{(n+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})}+2\left(\upsilon^{(n+1)}_{i}\right)^{2}+2\sigma_{i}\rho_{i}^{(n+1)}-4\sigma_{i}\left(\upsilon^{(n+1)}_{i}\right)\right)}}{{\left(2\left(\upsilon^{(n+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})}\right)}},$ (27)
@@
$\boldsymbol{f}_{i,j}^{(n+1)}(m)$ $\triangleq{\xi_{i}}\bigg{/}{\left(\upsilon^{(n+1)}_{i}\sqrt{V_{i}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})}\right)}\left(\boldsymbol{w}^{(n+1)}_{j}\right)^{H}\boldsymbol{h}_{i}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{l}_{i}\Lambda_{m}{\boldsymbol{L}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(n+1)}_{j},m\in M,j\in K,j\neq k,$ (28)
@@
$x_{k}^{(n+1)}$ $\triangleq{x}_{1,k}^{(n+1)}+{x}_{1,k_{e}}^{(n+1)}+x_{2,k}^{(n+1)}+x_{2,k_{e}}^{(n+1)}+z_{1,ke}^{(n+1)}-\left(e^{j\boldsymbol{\theta}^{(n)}}\right)^{H}\boldsymbol{\varphi}_{k}^{(n+1)}e^{j\boldsymbol{\theta}}-2\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(n+1)}\right)M,$ (30)
@@
$\boldsymbol{y}^{(n+1)}_{k}(m)$ $\triangleq\boldsymbol{y}^{(n+1)}_{k,k}(m)+\boldsymbol{y}^{(n+1)}_{k_{e}}(m)+\boldsymbol{f}^{(n+1)}_{k}(m)+\boldsymbol{f}^{(n+1)}_{k_{e}}(m)+{\sum}_{m=1}^{N}e^{-j\theta_{m}^{(n)}}\boldsymbol{\varphi}_{k}^{(n+1)}(m,n)+\lambda_{\text{max}}\left(\boldsymbol{\varphi}_{k}^{(n+1)}\right).$ (31)